# Differentially Private Low-Rank Adaptation of Large Language Model Using Federated Learning

### v1

@@ 5.1. Tasks and Datasets
To mimic the specific implementation scenario, we chose three datasets to test our algorithm. To mimic the protection of highly sensitive data, we choose the financial and medical data which always contains sensitive information. We also test our algorithm on a normal dataset to provide a comparison with other methods.
@@ 5.2. Experiments Setup
Experiments settings: For federated learning, We split the dataset evenly and stored it locally among all $K$ nodes. We adopt an average for aggregating local model weights and we set all $\rho_{k}$ in Equation 4 to $\frac{1}{K}$. We assume that the communication channels are noiseless and ideal. Each node $k$ uploads its $\bm{\theta}_{k}$ to the server in every iteration. We choose $K$ as $5$ and we set the maximum iteration $T$ as 50. For the training, we follow the settings in 2.
@@ 5.3. Results on Medical Tasks
In our experiments, we evaluated the performance of several models in the context of differential privacy for medical tasks. We investigated the impact of varying the parameters $\epsilon$ and $\delta$ for the language model fine-tuning on the medical datasets. In all experiments, the compression rank was fixed at $512$ and the $\delta$ is set to $51e-$ 5 when we varying the $\epsilon$ and the $\epsilon$ is settd to $2$.
One notable trend is the general decrease in performance across all models with stricter privacy settings (lower $\epsilon$ and higher $\delta$ values). For instance, GPT-2’s performance on MedQuAD drops from 69.2 at the original setting to 55.3 and 49.1 when $\epsilon$ is reduced to 2 and $\delta$ to $051e-$ 05, respectively (Tables 3 and 7). This trend indicates a trade-off between privacy and utilities.
Table 3. The performance of the global model under different $\epsilon$ for GPT-2 in general and Medical tasks.
$\epsilon$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
origin 61.9 53.8 48.3 69.2 39.5 55.1
2 51.5 46.3 46.1 55.3 27.0 49.1
4 49.0 45.4 46.7 57.3 37.1 47.4
6 58.4 41.5 45.6 63.5 38.9 52.9
8 52.9 45.1 42.7 66.6 39.1 47.9
10 58.9 53.1 44.7 64.1 35.9 51.8
Table 4. The performance of the global model under different $\epsilon$ for Bert in general and Medical tasks.
$\epsilon$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
origin 45.8 51.3 41.2 69.1 56.4 73.3
2 36.2 41.0 29.8 53.2 46.8 66.5
4 33.1 37.2 38.0 63.6 46.3 61.0
6 40.1 40.2 33.1 68.8 47.6 60.9
8 40.1 42.2 38.9 60.3 51.5 70.2
10 42.2 45.9 37.9 68.9 51.6 69.2
Table 5. The performance of the global model under different $\epsilon$ for ChatGLM-6B in general and Medical tasks.
$\epsilon$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
origin 44.2 51.2 50.1 66.6 71.9 69.3
@@
Table 6. The performance of the global model under different $\epsilon$ for Llama-7B in general and Medical tasks.
$\epsilon$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
origin 71.3 66.2 65.5 73.1 69.4 65.5
@@
In contrast, some models like Llama-7B show a more resilient performance under varying $\epsilon$ values. For example, its performance on LiveQA only marginally decreases from 69.4 to 66.1 when $\epsilon$ is increased from the original to 10 (Table 6). This suggests that certain models might be better suited for privacy-sensitive applications.
Additionally, the impact of changing $\delta$ values appears to be more model-specific. Bert’s performance on MEDIQA-Ans decreases significantly from 73.3 in the original setting to 57.4 when $\delta$ is reduced to $061e-$ 06 (Table 8), highlighting a potentially higher sensitivity to $\delta$ adjustments.
Table 7. The performance of the global model under different $\delta$ for GPT-2 in general and Medical tasks.
$\delta$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
origin 61.9 53.8 48.3 69.2 39.5 55.1
1e-02 60.2 50.7 46.6 69.1 32.5 51.1
1e-03 54.7 49.5 38.7 59.8 36.1 46.1
1e-04 61.6 47.6 35.0 68.5 30.9 50.3
1e-05 51.5 46.3 46.1 55.3 27.0 49.1
1e-06 53.4 42.9 40.6 69.1 30.5 41.4
Table 8. The performance of the global model under different $\delta$ for Bert in general and Medical tasks.
$\delta$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
origin 45.8 51.3 41.2 69.1 56.4 73.3
1e-02 37.7 44.3 37.8 65.3 50.9 73.3
1e-03 37.3 44.1 38.6 65.4 48.2 67.5
1e-04 35.1 40.1 33.6 64.3 52.5 59.1
1e-05 36.2 41.0 29.8 53.2 46.8 66.5
1e-06 31.3 43.2 36.4 62.9 40.6 57.4
Table 9. The performance of the global model under different $\delta$ for ChatGLM-6B in general and Medical tasks. The left column denotes the maximum value of $\delta$.
$\delta$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
origin 44.2 51.2 50.1 66.6 71.9 69.3
@@
Table 10. The performance of the global model under different $\delta$ for Llama-7B in general and Medical tasks. The left column denotes the maximum value of $\delta$.
$\delta$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
origin 71.3 66.2 65.5 73.1 69.4 65.5
@@
5.4. Results on Financial Tasks
In this section, we compare the financial tasks performance of the global model which is fine-tuned on the financial datasets. In our experiments, We assume that there are $5$ clients in participating the whole training process. We use the above-mentioned models as our initial models. All the local models and global models follow the same size and use the original model weights in the initial. We trained each local model under the same set of parameters, such as learning rate, decay strategy, and training epochs. We set the compression rank $r$ to 512 and $\epsilon$ value as 2 when we vary the $\delta$ and $\delta$ as $51e-$ 5 when we vary the $\epsilon$. The origin row means that we train our model without differential private adaptations.
Table 11. The performance of the global model under different $\delta$ for GPT-2 in general and financial tasks. The left column denotes the maximum value of $\delta$.
$\delta$ BoolQ PIQA WinoGrande FPB FiQA SA TFNS
origin 41.1 39.9 37.5 59.8 73.8 66,0
1e-02 36.4 39.6 35.7 57.1 71.4 57.0
1e-03 34.7 35.4 29.4 51.1 63.3 52.2
1e-04 33.1 28.6 27.8 42.1 57.5 45.1
1e-05 25.5 20.5 22.2 39.1 42.9 34.3
1e-06 23.2 19.0 20.3 28.1 52.2 34.0
Table 12. The performance of the global model under different $\delta$ for Bert in general and financial tasks. The left column denotes the maximum value of $\delta$.
$\delta$ BoolQ PIQA WinoGrande FPB FiQA SA TFNS
origin 45.3 47.5 51.3 65.5 60.1 57.3
1e-02 41.8 41.7 46.7 59.6 59.1 55.6
1e-03 37.9 39.7 44.9 50.6 55.6 50.9
1e-04 37.7 36.8 39.2 48.4 51.4 47.9
1e-05 21.7 29.9 38.8 43.7 47.5 44.1
1e-06 20.3 28.7 30.2 37.3 45.3 40.3
Table 13. The performance of the global model under different $\delta$ for ChatGLM-6B in general and financial tasks. The left column denotes the maximum value of $\delta$.
$\delta$ BoolQ PIQA WinoGrande FPB FiQA SA TFNS
origin 61.5 69.3 63.8 55.5 74.3 70.1
1e-02 55.5 68.7 61.1 50.7 69.4 62.2
1e-03 53.6 65.3 54.9 49.1 64.5 56.2
1e-04 50.7 62.3 51.7 47.4 58.9 47.6
1e-05 44.9 58.3 48.6 29.9 48.1 42.5
1e-06 38.5 55.2 47.4 30.6 51.5 38.0
Table 14. The performance of the global model under different $\delta$ for Llama-7B in general and financial tasks. The left column denotes the maximum value of $\delta$.
$\delta$ BoolQ PIQA WinoGrande FPB FiQA SA TFNS
origin 76.3 79.7 70.0 59.1 80.0 69.6
1e-02 88.1 80.6 69.3 80.3 81.9 67.3
1e-03 81.8 78.2 57.6 64.1 75.0 65.4
1e-04 75.4 69.3 59.1 49.1 76.1 70.3
1e-05 59.3 71.5 61.6 38.5 48.3 43.2
1e-06 69.5 56.6 63.5 35.2 51.5 28.9
Table 15. The performance of the global model under different $\epsilon$ for GPT-2 in general and financial tasks. The left column denotes the maximum value of $\epsilon$.
$\epsilon$ BoolQ PIQA WinoGrande FPB FiQA SA TFNS
origin 41.1 39.9 37.5 59.8 73.8 66
2 25.5 20.5 22.2 39.1 42.9 34.3
4 26.8 19.5 29.3 40.8 53.9 35.2
6 37.3 26.5 28.2 49.0 55.5 49.6
8 33.9 27.5 31.1 53.2 56.9 54.7
10 39.5 31.2 31.2 53.9 57.1 54.8
Table 16. The performance of the global model under different $\epsilon$ for Bert in general and financial tasks. The left column denotes the maximum value of $\epsilon$.
$\epsilon$ BoolQ PIQA WinoGrande FPB FiQA SA TFNS
origin 45.3 47.5 51.3 65.5 60.1 57.3
2 21.7 29.9 38.8 43.7 47.5 44.1
4 24.5 33.5 39.9 45.2 47.3 44.5
6 27.5 34.6 45.0 49.9 53.8 45.7
8 31.0 36.2 46.7 50.4 54.9 51.9
10 29.1 37.6 47.5 52.3 59.9 49.2
Table 17. The performance of the global model under different $\epsilon$ for ChatGLM-6B in general and financial tasks.. The left column denotes the maximum value of $\epsilon$.
$\epsilon$ BoolQ PIQA WinoGrande FPB FiQA SA TFNS
origin 61.5 69.3 63.8 55.5 74.3 70.1
2 44.9 58.3 48.6 29.9 48.1 42.5
4 43.2 58.3 48.3 40.2 57.9 54.8
6 48.4 59.1 51.4 43.5 62.2 56.3
8 52.5 60.9 58.0 45.6 68.1 63.2
10 55.3 67.0 62.2 53.2 72.6 64.9
Table 18. The performance of the global model under different $\epsilon$ for Llama-7B in general and financial tasks. The left column denotes the maximum value of $\epsilon$.
$\epsilon$ BoolQ PIQA WinoGrande FPB FiQA SA TFNS
origin 76.3 79.7 70.0 59.1 80.0 69.6
2 59.6 71.5 61.6 38.5 48.3 43.2
4 63.4 69.3 61.9 43.1 65.0 65.4
6 71.9 70.3 59.1 49.1 76.1 65.3
8 76.4 80.4 69.9 68.5 73.1 77.2
10 87.9 80.1 71.1 69.9 81.5 82.9
The experimental data consistently demonstrate a clear trade-off between the privacy level (as adjusted by $\epsilon$ and $\delta$ values) and the model performance across various tasks. While all models show a decrease in performance with stricter privacy settings, the extent of this decrease varies, suggesting differences in how each model adapts to privacy constraints.
One notable trend is the general decrease in performance across all models with stricter privacy settings (lower $\epsilon$ and $\delta$ values). For instance, GPT-2’s performance on MedQuAD drops from 69.2 at the original setting to 55.3 and 49.1 when $\epsilon$ is reduced to 2 and $\delta$ to $051e-$ 05, respectively (Tables 3 and 7). This trend indicates a trade-off between privacy and effectiveness.
In contrast, some models like Llama-7B show a more resilient performance under varying $\epsilon$ values. For example, its performance on LiveQA only marginally decreases from 69.4 to 66.1 when $\epsilon$ is increased from the original to 10 (Table 6). This suggests that certain models might be better suited for privacy-sensitive applications.
Additionally, the impact of changing $\delta$ values appears to be more model-specific. Bert’s performance on MEDIQA-Ans decreases significantly from 73.3 in the original setting to 57.4 when $\delta$ is reduced to $061e-$ 06 (Table 8), highlighting a potentially higher sensitivity to $\delta$ adjustments.
Table 19. We show the communication overhead of GPT-2 in federated learning. We use the model parameter amount to denote the communication efficiency and compression ratio as an improvement indicator for different sets of rank $r$. The origin row means the results without our algorithm. For all the experiments, we train our local model under privacy parameter $\epsilon=8$ and $\delta=10e^{-5}$. We also include the model performance on general tasks.
rank $r$ BoolQ PIQA WinoGrande Communication overhead Reduction ratio
origin 61.9 56.9 48.9 127.4M -
512 39.4 41.7 45.5 101.3M 79.5%
256 38.2 44.3 43.2 82.1M 64.4%
128 32.1 50.3 45.9 69.5M 54.6%
64 32.4 49.1 39.9 58.6M 46.0%
32 33.5 46.3 39.4 58.0M 45.5%
Table 20. We show the communication overhead of Bert in federated learning. We use the model parameter amount to denote the communication efficiency and compression ratio as an improvement indicator for different sets of rank $r$. The origin row means the results without our algorithm. For all the experiments, we train our local model under privacy parameter $\epsilon=8$ and $\delta=10e^{-5}$. We also include the model performance on general tasks.
rank $r$ BoolQ PIQA WinoGrande Communication overhead Reduction ratio
origin 49.8 55.5 60.2 110.8M -
512 35.5 44.7 41.3 83.7M 75.5%
256 34.2 47.3 40.3 69.9M 63.1%
128 35.1 46.2 43.8 54.5M 49.2%
64 33.2 39.1 38.9 48.1M 43.4%
32 33.7 37.0 37.8 46.6M 42.0%
Table 21. We show the communication overhead of Llama-7B in federated learning. We use the model parameter amount to denote the communication efficiency and compression ratio as an improvement indicator for different sets of rank $r$. The origin row means the results without our algorithm. For all the experiments, we train our local model under privacy parameter $\epsilon=8$ and $\delta=10e^{-5}$. We also include the model performance on general tasks.
rank $r$ BoolQ PIQA WinoGrande Communication overhead Reduction ratio
origin 76.3 79.7 70.0 6.7B -
1024 78.8 83.1 71.1 2.43B 36.27%
@@ 5.4. Results on Financial Tasks
Table 22. We show the communication overhead of ChatGLM-6B in federated learning. We use the model parameter amount to denote the communication efficiency and compression ratio as an improvement indicator for different sets of rank $r$. The origin row means the results without our algorithm. For all the experiments, we train our local model under privacy parameter $\epsilon=8$ and $\delta=10e^{-5}$. We also include the model performance on general tasks.
rank $r$ BoolQ PIQA WinoGrande Communication overhead Reduction ratio
origin 69.5 76.3 71.3 6.2B -
1024 57.3 65.3 55.5 2.94B 47.4%
@@ 5.5. Reduction in Communication Overhead
GPT-2 (Table 19) showed a performance drop in BoolQ from 61.9 to 33.5, PIQA from 56.9 to 46.3, and WinoGrande from 48.9 to 39.4 as the rank reduced from origin to 32, alongside an increase in compression ratio from 45.5 to 79.5. BERT (Table 20) exhibited a similar performance decline, for example, in WinoGrande from 60.2 to 37.8, with a compression improvement from 42.0 to 75.5. Llama-7B (Table 21), a larger model, maintained a relatively stable performance (e.g., BoolQ at 76.3 to 39.1, WinoGrande 70.0 to 67.8 ) despite significant parameter reduction (compression ratio from 7.31 to 36.27). In contrast, ChatGLM-6B (Table 22) showed a more pronounced performance degradation (BoolQ dropping from 69.5 to 50.1, WinoGrande from 71.3 to 46.3) with a notable decrease in parameters (compression ratio moving from 7.6 to 47.4). These results highlight the variability in how different models respond to parameter reduction and compression, underscoring the challenge of balancing communication efficiency with performance in federated learning environments.
@@ 6. Conclusion
In this paper, we have introduced DP-LoRA, an innovative approach to preserve data privacy in the low-rank adaptation of LLMs. DP-LoRA ensures differential privacy while significantly reducing communication overhead. By introducing noise into gradients to achieve ($\epsilon$, $\delta$)-differential privacy, we enhance communication efficiency. Empirical validation through experiments demonstrates the efficacy of our approach, rendering LLMs more secure for domain-specific fine-tuning. Looking ahead, our future research endeavors will explore refining the error bounds for differential privacy, particularly in the context of fine-tuning LLMs where the existing bounds may be less stringent. Additionally, we aim to broaden our experimental scope to encompass a wider array of domain-specific datasets, further validating the applicability and robustness of DP-LoRA.

### v2

@@ 5.1. Tasks and Datasets
To mimic the specific implementation scenario, we chose three datasets to test our algorithm. To mimic the protection of highly sensitive data, we choose the financial and medical data which always contains sensitive information. We also test our algorithm on a normal dataset to provide a comparison with other methods. In the following, we primarily present the results on the medical datasets, with the financial datasets provided in the appendix.
@@
Note that the aforementioned datasets often lack truly sensitive information due to their open-source nature. Our algorithm does not differentiate between sensitive and non-sensitive portions of the data; instead, it focuses on protecting the local dataset as a whole. Users can readily apply our algorithm to data they deem sensitive.
@@ 5.2. Experiments Setup
Experiments settings: For comparison, we choose three federated learning algorithms as our baseline. (1)Vanilla FL (2)T-FedAvg (Xu et al., 2020) (3)CMFL (Luping et al., 2019). We train the baselines and the proposed method in the same settings. We split the dataset evenly and stored it locally among all $K$ nodes. We adopt an average for aggregating local model weights and we set all $\rho_{k}$ in Equation 4 to $\frac{1}{K}$. We assume that the communication channels are noiseless and ideal. Each node $k$ uploads its $\bm{\theta}_{k}$ to the server in every iteration. We choose $K$ as $5$ and we set the maximum iteration $T$ as 50. For the training, we follow the settings in Table. 2.
@@ 5.3. Results on Medical Tasks
In our experiments, we evaluated the performance of several models in the context of differential privacy for medical tasks. We investigated the impact of varying the parameters $\epsilon$ and $\delta$ for the language model fine-tuning on the medical datasets. In all experiments, the compression rank was fixed at $512$ and the $\delta$ is set to $51e-$ 5 when we varying the $\epsilon$ and the $\epsilon$ is set to $2$.
One notable trend is the general decrease in performance across all models with stricter privacy settings (lower $\epsilon$ and lower $\delta$ values). For instance, Llama-7B’s performance on LiveQA drops from $69.4$ at the original setting to $55.9$ and $49.3$ when $\epsilon$ is reduced to $2$ and $\delta$ to $051e-$ 05, respectively (Table 3 and Table 4). This trend indicates a trade-off between privacy and utilities.
Table 3. The performance of the global model under different $\epsilon$ for ChatGLM-6B and Llama-7B in general and Medical tasks.
Model $\epsilon$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
ChatGLM-6B origin 44.2 51.2 50.1 66.6 71.9 69.3
@@
Llama-7B origin 71.3 66.2 65.5 73.1 69.4 65.5
@@
In contrast, some models like ChatGLM-6B show a more resilient performance under varying $\epsilon$ values. For example, its performance on LiveQA only marginally decreases from $71.9$ to $69.8$ when $\epsilon$ is increased from the original to 10 (Table 3). This suggests that certain models might be better suited for privacy-sensitive applications.
Table 4. The performance of the global model under different $\delta$ for ChatGLM-6B and Llama-7B in general and Medical tasks. The left column denotes the maximum value of $\delta$.
Model $\delta$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
ChatGLM-6B origin 44.2 51.2 50.1 66.6 71.9 69.3
@@
Llama-7B origin 71.3 66.2 65.5 73.1 69.4 65.5
@@
Table 5. We show the communication overhead of ChatGLM-6B and Llama-7B in federated learning. We use the model parameter amount to denote the communication efficiency and compression ratio as an improvement indicator for different sets of rank $r$. The origin row means the results without our algorithm. For all the experiments, we train our local model under privacy parameter $\epsilon=8$ and $\delta=10e^{-5}$. We also include the model performance on general tasks.
Model Method rank $r$ BoolQ PIQA WinoGrande Communication overhead Reduction ratio
Llama-7B Vanila FL (McMahan et al., 2017) - 76.3 79.7 70.0 6.7B -
CMFL (Luping et al., 2019) - 71.4 77.3 65.8 1.79B 26.8%
T-FedAvg (Xu et al., 2020) - 69.9 70.1 67.0 1.18B 17.7%
Ours 1024 78.8 83.1 71.1 2.43B 36.27%
@@
16 38.9 61.3 45.7 0.31B 4.63%
4 36.8 50.7 32.3 0.17B 2.53%
ChatGLM-6B Vanila FL (McMahan et al., 2017) - 69.5 76.3 71.3 6.2B -
CMFL (Luping et al., 2019) - 55.2 64.9 65.3 1.95B 31.4%
T-FedAvg (Xu et al., 2020) - 53.9 68.6 63.4 1.41B 22.9%
Ours 1024 57.3 65.3 55.5 2.94B 47.4%
@@ 5.4. Reduction in Communication Overhead
Llama-7B (Table 5), one of the most popular large language models, maintained a relatively stable performance (e.g., BoolQ at 76.3 to 39.1, WinoGrande 70.0 to 67.8 ) despite significant parameter reduction (compression ratio from 7.31 to 36.27). In contrast, ChatGLM-6B (Table 5) showed a more pronounced performance degradation (BoolQ dropping from 69.5 to 50.1, WinoGrande from 71.3 to 46.3) with a notable decrease in parameters (compression ratio moving from 7.6 to 47.4). These results highlight the variability in how different models respond to parameter reduction and compression, underscoring the challenge of balancing communication efficiency with performance in federated learning environments.
@@ 6. Conclusion
In this paper, we have introduced DP-LoRA, an innovative approach to preserve data privacy in the low-rank adaptation of LLMs. DP-LoRA ensures differential privacy while significantly reducing communication overhead. By introducing noise into gradients to achieve ($\epsilon$, $\delta$)-differential privacy, we enhance communication efficiency. Empirical validation through experiments demonstrates the efficacy of our approach, rendering LLMs more secure for domain-specific fine-tuning. Looking ahead, our future research endeavors will explore refining the error bounds for differential privacy, particularly in the context of fine-tuning LLMs where the existing bounds may be less stringent. Additionally, we aim to broaden our experimental scope to encompass a wider array of domain-specific datasets, further validating the applicability and robustness of DP-LoRA. Finally, we plan to conduct more experimental analysis to address the scenarios where different institutions possess their own unique datasets that do not conform to the same distribution as others.